# Hydrodynamic-like behaviour of glasma

### v1

@@ 1 Introduction
The authors of the study [9] found that as long as a system is boost-invariant, and its energy-momentum tensor is traceless and initially diagonal, the system at the early stage of its evolution behaves in some respects as a fluid. A similar finding was presented in [10, 11, 12], where the glasma from the earliest stage of ultrarelativistic heavy-ion collisions was studied. Although the system is far from equilibrium and its dynamics is governed by the Yang-Mills equations, glasma temporal evolution resembles that of hydrodynamics. Specifically, an approximate eccentricity scaling of the glasma elliptic flow was found.
@@
Throughout the paper we use the natural system of units with $c=\hbar=k_{B}=1$. Greek letters $\mu,\nu,\rho,\dots$ label components of four-vectors. Vectors transverse to the collision axis $z$ are denoted $\vec{x}_{\perp}$ and their components are indexed with Latin letters $i,j,k,\dots$. Latin letters from the beginning of the alphabet $a,b,c\dots$ label colour components of elements of the SU($N_{c}$) gauge group in the adjoint representation. We use three systems of coordinates: Minkowski $(t,z,x,y)$, light-cone $(+,-,x,y)$ and Milne $(\tau,\eta,x,y)$, where $x^{\pm}=(t\pm z)/\sqrt{2}$, $\tau=\sqrt{t^{2}-z^{2}}$ and $\eta=\ln(x^{+}/x^{-})/2$. The indices of vector and tensor components, like $A^{t}$, $F^{+x}$, $T^{\tau\eta}$, clearly show which coordinates are used.
@@ 3 Equation of universal flow
In this section we explain why the glasma is expected to behave similarly to a fluid. Our arguments are inspired by the study [9] where an equation for the flow transverse to the collision axis was derived adopting a few assumptions relevant for the early stage of ultrarelativistic-heavy-ion collisions. We arrive at the same equation, which we call ‘the equation of universal flow’, but our derivation presented in Sec. 3.1 is somewhat different. In Sec. 3.2 we explicitly show that the assumptions used to derive the equation of universal flow are satisfied by the energy-momentum obtained in [10] using an expansion in powers of $\tau$. We further show that the equation is exactly satisfied order by order by the energy-momentum tensor obtained with a proper time expansion up to seventh order.
@@ 3.1 Derivation
In Refs. [10, 11] we obtained the energy-momentum tensor of the glasma from Pb-Pb collisions using a proper time expansion of chromodynamic potentials. In Fig. 1 we show the left and right sides of Eq. (3.11) as contour plots in the transverse $x\!-\!y$ plane, for $t=0.06$ fm, $\eta=0$ and impact parameter $b=2$ fm. Fig. 1 clearly shows that Eq. (3.11) is satisfied with good accuracy.
@@ 4 Glasma as ideal or anisotropic fluid
Since the glasma initial longitudinal pressure is negative, we expect that ansiotropic hydrodynamics [6, 7] will provide a better description. We therefore consider
@@
Although a dissipative contribution is not explicitly included in Eq. (4.2), the corresponding hydrodynamics [6, 7] is not ideal. Entropy production proportional to ${\cal E}_{T}-{\cal E}_{L}$ is postulated, which provides an extra equation that is used to close the system of hydrodynamic equations. When ${\cal P}_{T}$ and ${\cal P}_{L}$ become equal to each other, the entropy production ceases and anisotropic hydrodynamics changes into the ideal version.
@@
We have applied the procedure described above to construct $T^{\mu\nu}_{\rm ideal}$ and $T^{\mu\nu}_{\rm aniso}$ from the glasma energy-momentum tensor $T^{\mu\nu}_{\rm glasma}$ computed using the methods developed in [10, 11, 23]. We displace the centers of two equal size colliding ions symmetrically about the $x$-axis. The energy-momentum tensor is therefore symmetric and has cylindrical symmetry. This means it is fully determined by the seven components $T^{tt},T^{tz},T^{tx},T^{zz},T^{zx},T^{xx},T^{xy}$. For zero impact parameter $T^{tz}$ and $T^{tx}$ are identically zero.
@@
The elliptic flow observed in relativistic heavy-ion collisions is zero or almost zero for $b=0$ and grows as $b$ increases [29]. The system’s initial eccentricity, $\epsilon(0)$, behaves similarly and consequently the ratio $v_{2}/\epsilon(0)$ is independent or weakly dependent on impact parameter. This is called eccentricity scaling and is generally taken as an indicator of fluid behaviour [1, 2]. The idea is that in hydrodynamics the difference of pressure gradients in and out the reaction plane, which is encoded in the system’s spatial eccentricity, is responsible for the azimuthal asymmetry of the transverse collective flow quantified by $v_{2}$.
It was demonstrated in [10, 11, 12] that the ratio of the glasma elliptic flow $v_{2}(t)$ at short but finite time $t$ and the initial eccentricity $\epsilon(0)$ shows weak eccentricity scaling. The ratio $v_{2}(t)/\epsilon(0)$ has a dependence on $b$ which is noticeably weaker than that of $v_{2}(t)$ and $\epsilon(0)$. Specifically, when the impact parameter grows from 1 to 6 fm, the elliptic coefficients $v_{2}(t)$ at $t=0.06$ fm increases by the factor of 10 but the ratio $v_{2}(t)/\epsilon(0)$ decreases only by a factor of 3. This is shown in Fig. 5. The quantities $v_{2}(t)$ and $\epsilon(0)$ are calculated in three different ways: directly from the glasma energy-momentum tensor and from the energy-momentum tensors of ideal and anisotropic hydrodynamics. Both ideal and anisotropic hydrodynamics give results that are very similar to what is obtained from the glasma. This shows that glasma transverse dynamics is well represented by hydrodynamics.

### v3

@@ 1 Introduction
The authors of the study [9] found that as long as a system is boost-invariant, and its energy-momentum tensor is traceless and initially diagonal, the system at the early stage of its evolution behaves in some respects as a fluid. A similar finding was presented in [10, 11, 12], see also Sec. 4.5.2 of Ref. [13], where the glasma from the earliest stage of ultrarelativistic heavy-ion collisions was studied. Although the system is far from equilibrium and its dynamics is governed by the Yang-Mills equations, glasma temporal evolution resembles that of hydrodynamics. Specifically, an approximate eccentricity scaling of the glasma elliptic flow was found [10, 11, 12].
@@
Throughout the paper we use the natural system of units with $c=\hbar=k_{B}=1$. Greek letters $\mu,\nu,\rho,\dots$ label components of four-vectors. Vectors transverse to the collision axis $z$ are denoted $\vec{x}_{\perp}$ and their components are indexed with Latin letters $i,j,k,\dots$. Latin letters from the beginning of the alphabet $a,b,c\dots$ label colour components of elements of the SU($N_{c}$) gauge group in the adjoint representation. We use three systems of coordinates: Minkowski $(t,z,x,y)$, light-cone $(x^{+},x^{-},x,y)$ and Milne $(\tau,\eta,x,y)$, where $x^{\pm}=(t\pm z)/\sqrt{2}$, $\tau=\sqrt{t^{2}-z^{2}}$ and $\eta=\ln(x^{+}/x^{-})/2$. The indices of vector and tensor components, like $A^{t}$, $F^{+x}$, $T^{\tau\eta}$, clearly show which coordinates are used.
@@ 3 Equation of universal flow
In this section we explain why the glasma is expected to behave similarly to a fluid. Our arguments are inspired by the study [9] where an equation for the flow transverse to the collision axis was derived adopting a few assumptions relevant for the early stage of ultrarelativistic-heavy-ion collisions. We arrive at the same equation, which we call ‘the equation of universal flow’, but our derivation presented in Sec. 3.1 is somewhat different. In Sec. 3.2 we explicitly show that the assumptions used to derive the equation of universal flow are satisfied by the energy-momentum tensor obtained in [10] using an expansion in powers of $\tau$. A similar analysis is given in Sec. 4.5.2 of Ref. [13]. We further show that the equation is exactly satisfied order by order by the energy-momentum tensor obtained with a proper time expansion up to seventh order.
@@ 3.1 Derivation
In Refs. [10, 11] we obtained the energy-momentum tensor of the glasma from Pb-Pb collisions using a proper time expansion of chromodynamic potentials. In Fig. 1 we show the left and right sides of Eq. (3.11) as contour plots in the transverse $x$-$y$ plane, for $t=0.06$ fm, $\eta=0$ and impact parameter $b=2$ fm. Fig. 1 clearly shows that Eq. (3.11) is satisfied with good accuracy. To obtain a quantitative measure we calculate
$\Delta=\frac{|L|-|R|}{|L|+|R|}=0.184$ (3.12)
where $|L|$ and $|R|$ indicate the average over the transverse plane of the absolute value of the left and right sides of equation (3.11).
@@ 4 Glasma as ideal or anisotropic fluid
Since the glasma initial longitudinal pressure is negative, we expect that anisotropic hydrodynamics [6, 7] will provide a better description. We therefore consider
@@
Although a dissipative contribution is not explicitly included in Eq. (4.2), the corresponding hydrodynamics [6, 7] is not ideal. Entropy production proportional to ${\cal P}_{T}-{\cal P}_{L}$ is postulated, which provides an extra equation that is used to close the system of hydrodynamic equations. When ${\cal P}_{T}$ and ${\cal P}_{L}$ become equal to each other, the entropy production ceases and anisotropic hydrodynamics changes into the ideal version.
@@
We have applied the procedure described above to construct $T^{\mu\nu}_{\rm ideal}$ and $T^{\mu\nu}_{\rm aniso}$ from the glasma energy-momentum tensor $T^{\mu\nu}_{\rm glasma}$ computed using the methods developed in [10, 11, 24]. We displace the centers of two equal size colliding ions symmetrically about the $x$-axis. The energy-momentum tensor is symmetric which means it is fully determined by the ten components: $T^{tt},T^{tz},T^{tx},T^{ty},T^{zz},T^{zx},T^{zy},T^{xx},T^{xy},T^{yy}$. At $z=0$, the components $T^{tz}$, $T^{zx}$ and $T^{zy}$ are identically zero. For zero impact parameter there is cylindrical symmetry and thus $T^{tx}=T^{ty}$, $T^{zx}=T^{zy}$ and $T^{xx}=T^{yy}$. Consequently, when $b=0$ and $z=0$ the energy-momentum tensor is determined by the 5 non-zero components.
@@
The elliptic flow observed in relativistic heavy-ion collisions is zero or almost zero for $b=0$ and grows as $b$ increases [30]. The system’s initial eccentricity, $\epsilon(0)$, behaves similarly and consequently the ratio $v_{2}/\epsilon(0)$ is independent or weakly dependent on impact parameter [31]. This is called eccentricity scaling and is generally taken as an indicator of fluid behaviour [1, 2]. The idea is that in hydrodynamics the difference of pressure gradients in and out the reaction plane, which is encoded in the system’s spatial eccentricity, is responsible for the azimuthal asymmetry of the transverse collective flow quantified by $v_{2}$.
It was demonstrated in [10, 11, 12] that the glasma elliptic flow $v_{2}(t)$ coefficient at small but finite time $t$ and the initial eccentricity $\epsilon(0)$ behave in a similar way as functions of impact parameter. The ratio $v_{2}(t)/\epsilon(0)$ does not show eccentricity scaling but it has a dependence on $b$ which is noticeably weaker than that of $v_{2}(t)$ and $\epsilon(0)$. Specifically, when the impact parameter grows from 1 to 6 fm, the elliptic coefficients $v_{2}(t)$ at $t=0.06$ fm increases by the factor of 10 but the ratio $v_{2}(t)/\epsilon(0)$ decreases only by a factor of 3. This is shown in Fig. 5. The quantities $v_{2}(t)$ and $\epsilon(0)$ are calculated in three different ways: directly from the glasma energy-momentum tensor and from the energy-momentum tensors of ideal and anisotropic hydrodynamics. Both ideal and anisotropic hydrodynamics give results that are very similar to what is obtained from the glasma. This shows that glasma transverse dynamics can be approximated by hydrodynamics.
@@ 6 Summary and conclusions
Finally, we mention some possible improvements and future directions for this work. One important simplification in our approach is the assumption of boost invariance. Clearly it would be desirable to go beyond this approximation and include some of the effects of the longitudinal dynamics. Unfortunately this would be very difficult because it would require a modification of the ansatz that determines the structure of the gauge potential.
It would be also desirable to check whether the fluid-like behaviour of glasma produced in heavy-ion collisions also occurs in smaller systems obtained from proton-nucleus or proton-proton collisions. Our approach could be easily extrapolated to small mass numbers of colliding nuclei but the extent to which these results would be reliable is unclear. Since our method is fully classical, we assume that the system’s dynamics is dominated by long-wavelength modes. This assumption becomes less reliable for smaller systems because the role of short-wavelength quantum modes increases.